Show via probability manipulation that the following is true: P(A∩B∩C) = P(A|B∩C) x P(B|C) x P(C)
Answer #1

Solution:

Conditional probability rule shows that

P(A|B)= P(A∩B)/P(B)

P(A ∩ B) = P(A | B)P(B) = P(B | A)P(A)......multiplication rule.

So, P(A∩(B∩C))= P(A|(B∩C))* P(B∩C)..........(i)

In the same manner by the multiplication rule:

P(B∩C) = P(B|C) x P(C).....(ii)

Putting the (ii) in to (i), we get

P(A∩B∩C) = P(A|B∩C) x P(B|C) x P(C)